OpenStudy (anonymous):

HELP PLEASE! I NEED ANSWERS! Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = -1.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

you know what this looks like?

OpenStudy (anonymous):



OpenStudy (anonymous):

vertex is half way between the focus and the directrix, namely at \((-5,2)\)

OpenStudy (anonymous):

that means it looks like \[4p(y-2)=(x+5)^2\]and all you need is \(p\) the distance between the focus and the vertex, which is 3

OpenStudy (anonymous):

final answer \[12(y-2)=(x+5)^2\]

OpenStudy (anonymous):

check the answer here http://www.wolframalpha.com/input/?i=parabola+12%28y-2%29%3D%28x%2B5%29^2